OpenStudy (anonymous):

Solve the system using elimination. {2x+6y=-12 {5x-5y=10

OpenStudy (owlcoffee):

You say we have: {2x+6y=-12 {5x-5y=10 Okay, what we need to do, in order to elimiate either x or y, would be multiplying any of those two ecuations by a number, so we have two opposites. I'll multiply the first one by -5 and the second one by 2. We'll end up having: {-10x-30y=-60 {10x-10y=20 So we can simplify the Xs and sum the ecuations with only y on them. {-30y=-60 {-10y=20 We'll end up having: -40y=-40 => y=-40/-40 => y=1 Now we know that y01, you just take one of the other initial ecuations and replace it. So, we have for example: 2x+6y=-12 ; y=1 => 2x+6(1)=-12 We'll end up having: 2x=-18 then: x= -18/2 x=-9. And all you have to do is veify if the solutions are correct :)